OpenStudy (anonymous):

Find the common ratio for the geometric series 1/2,-1/4,1/8,-1/16....

OpenStudy (anonymous):

Formula r= t1/tn-1 tn=t1r^n-1

OpenStudy (anonymous):

\[ratio=\frac{ 2^{nd} term }{ 1^{st} term } = \frac{ 3^{rd} term }{ 2^{nd} term }\] and so on

OpenStudy (anonymous):

r =-1/2

OpenStudy (anonymous):

Thought t1=1/2

OpenStudy (anonymous):

-1/4 over 1/2 =-1/4x2 = -1/2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

WHat about 1/8

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ 8 } }{ -\frac{ 1 }{ 4 } }\] will give the same result as \[\frac{ -\frac{ 1 }{ 4 } }{ \frac{ 1 }{ 2 } }\]

OpenStudy (anonymous):

ok cool give this guy a medal please @RadEn

OpenStudy (anonymous):

So -1/16/1/8